A bag of coins contains pennies, dimes, and quarters (and no other types of coin or currency). There are twice as many dimes as pennies, and there are three times as many quarters as dimes. Which of the following could be the amount of money in the bag?
Answer by richwmiller(17219)

so what are the choices or should we guess?
so far we know
d=2p
q=3d
we have
p+10d+25q=n
p+10*2p+25*3*2p=n
p+20p+150p=n
171p=n
d=2p
q=3*2p=6p